solve the following system of equations using matrices
x-y+4z=6
2x+z=1
x+5y+z=-9
is it (0, 1, -2)??

Answer by funmath(2873)

solve the following system of equations using matrices
x-y+4z=6
2x+z=1
x+5y+z=-9
is it (0, 1, -2)??
:
Almost, it's (0,-2,1), since you have all the right numbers but they're out of order. Check to make sure that your second equation is 2x+0y+z=1. Also, check to make sure that you didn't simply write the z and y values in the wrong position when you wrote in your answer. The order is alphabetical (x,y,z). Let me know if you want me to solve it for you and what method you nee to use.
Happy Calculating!!!